# A Classification Benchmark for Artificial Intelligence Detection of Laryngeal Cancer from Patient Speech

### v1

@@ 3.3.1 Predictive Performance
In addition to the above metrics, we also suggest the area under the receiver operating characteristic curve (AUROC). The ROC curve plots the true positive rate against the false positive rate at different decision thresholds. AUROC is particularly useful for evaluating models when the positive class is small compared to the negative class, which is common in the datasets used in this area.
@@ 3.3.2 Fairness
In medical application, it is important to assess the fairness of the algorithms being developed. As such we suggest conductng statistical tests to evaluate the relationship between classifier performance and patient demographics. The most commonly available patient demographics are gender and age of patients. For gender, we propose using a Fisher’s Exact Test to examine any association between classifier performance and patient gender (Sprent, 2011). We propose a Fisher’s Exact test as it is used for categorical variables and widely recommended for small sample sizes (Kim, 2016). For age, we suggest a t-test to compare performance across different age groups (Kalpić et al., 2011). Supplementary figures may also be helpful and should be included in supplementary material or code repositories. These figures should include more details on the correctly and incorrectly classified patients, such as stacked bar charts or boxplots.
@@ 4 Results and Discussion
All experiments were performed using Python 3.10.12 on a machine running Windows 11 with an Intel Core i7-1260P CPU with 16GB RAM. All code and models are available on github at https://github.com/mary-paterson/LaryngealCancerClassificationBenchmark
@@ 4.1 Predictive Performance
Table 4 shows the results of the classifiers on both the FEMH (holdout) and SVD (external) test sets. The bolded numbers indicate the best value for each metric on the test set. The results show that adding demographics consistently improves classifier performance. This improvement is expected, as the demographic differences between classes in both datasets align with patterns likely to occur in real-world applications.
On the holdout test set, when using voice only, FeatureStates achieved the best performance for both the SVM and MLP classifiers, while MFCC performs the best for logistic regression. On the external test set, FeatureStates outperformed all other features across all three algorithms. This outcome aligns with expectations, since Wav2Vec2, the source of FeatureStates, was trained on multiple datasets and designed to be generalisable across different recording environments and devices, making it best suited for both holdout and external test sets.
When demographics are used alongside voice for classification, OpenSmile performs best across all three algorithms for the holdout test set. On the external test set, OpenSmile performed best for MLP and logistic regression, while MFCC was best for SVM. It can be seen in Figure 4 that the audio features used had more of an effect on performance than the algorithm used with OpenSmile consistently outperforming FeatureStates and MFCC when combined with patient demographics.
The classifiers perform worse on the external test data than the holdout data. This is expected due to several differences between the datasets, including recording devices, recording environments, recording lengths, and demographics. The two datasets were also collected in different countries, so the speakers’ accents may have an effect, although this is limited by using the elongated vowel task. The benign groups also include different pathologies. It is difficult to disentangle these differences to understand what the leading cause of the performance reduction is or whether all of these factors have an effect. Figure 4 shows the difference in balanced accuracy between the holdout (FEMH) and external (SVD) test sets. Although the performance on the external dataset is generally worse, the confidence intervals overlap for most classifiers except for MFCC when using voice only for logistic regression.
Figure 4: The balanced accuracy for the holdout (FEMH) and external (SVD) test sets. 95% confidence intervals are shown; notably, the confidence interval cannot be seen for OpenSmile logistic regression with voice and demographics as input; this is because all samples were classified as malignant.
{adjustwidth}
@@
-1.2cm-1.2cm FEMH (holdout test set) SVD (external test set) FeatureStates OpenSmile MFCC FeatureStates OpenSmile MFCC SVM Voice Balanced Accuracy 0.691 (0.593, 0.784) 0.680 (0.580, 0.772) 0.631 (0.531, 0.731) 0.628 (0.557, 0.696) 0.582 (0.497, 0.662) 0.532 (0.462, 0.613) Sensitivity 0.680 (0.480, 0.857) 0.640 (0.440, 0.818) 0.400 (0.206, 0.609) 0.763 (0.621, 0.900) 0.500 (0.341, 0.667) 0.289 (0.154, 0.455) Specificity 0.702 (0.665, 0.736) 0.720 (0.686, 0.754) 0.861 (0.835, 0.887) 0.493 (0.466, 0.521) 0.663 (0.637, 0.690) 0.775 (0.752, 0.799) AUC 0.760 (0.665, 0.845) 0.742 (0.676, 0.805) 0.701 (0.594, 0.807) 0.649 (0.562, 0.735) 0.615 (0.522, 0.711) 0.592 (0.503, 0.679) Voice + Demographics Balanced Accuracy 0.700 (0.594, 0.789) 0.752 (0.656, 0.837) 0.701 (0.599, 0.803) 0.609 (0.529, 0.691) 0.646 (0.566, 0.730) 0.681 (0.599, 0.759) Sensitivity 0.640 (0.429, 0.818) 0.760 (0.571, 0.923) 0.600 (0.400, 0.800) 0.526 (0.364, 0.688) 0.526 (0.370, 0.688) 0.605 (0.442, 0.756) Specificity 0.759 (0.726, 0.793) 0.743 (0.709, 0.776) 0.802 (0.770, 0.832) 0.691 (0.666, 0.717) 0.765 (0.741, 0.787) 0.758 (0.735, 0.781) AUC 0.773 (0.682, 0.855) 0.802 (0.743, 0.851) 0.807 (0.720, 0.881) 0.716 (0.642, 0.786) 0.772 (0.714, 0.829) 0.769 (0.699, 0.831) MLP Voice Balanced Accuracy 0.683 (0.576, 0.784) 0.632 (0.533, 0.727) 0.581 (0.478, 0.688) 0.628 (0.547, 0.705) 0.610 (0.529, 0.689) 0.548 (0.477, 0.616) Sensitivity 0.520 (0.318, 0.714) 0.560 (0.360, 0.750) 0.480 (0.280, 0.697) 0.526 (0.368, 0.676) 0.579 (0.417, 0.735) 0.763 (0.622, 0.889) Specificity 0.846 (0.817, 0.873) 0.704 (0.668, 0.740) 0.682 (0.646, 0.718) 0.729 (0.706, 0.753) 0.640 (0.615, 0.667) 0.333 (0.308, 0.358) AUC 0.754 (0.643, 0.857) 0.709 (0.637, 0.779) 0.638 (0.514, 0.746) 0.692 (0.607, 0.772) 0.633 (0.538, 0.727) 0.575 (0.489, 0.657) Voice + Demographics Balanced Accuracy 0.629 (0.522, 0.730) 0.755 (0.659, 0.841) 0.655 (0.556, 0.762) 0.669 (0.586, 0.748) 0.696 (0.612, 0.775) 0.657 (0.573, 0.742) Sensitivity 0.440 (0.227, 0.640) 0.760 (0.571, 0.923) 0.440 (0.250, 0.654) 0.553 (0.390, 0.704) 0.658 (0.488, 0.813) 0.500 (0.333, 0.667) Specificity 0.819 (0.786, 0.847) 0.750 (0.717, 0.784) 0.871 (0.844, 0.897) 0.786 (0.764, 0.808) 0.734 (0.708, 0.758) 0.815 (0.794, 0.835) AUC 0.758 (0.663, 0.833) 0.809 (0.753, 0.858) 0.805 (0.725, 0.881) 0.715 (0.627, 0.789) 0.758 (0.695, 0.822) 0.762 (0.695, 0.826) Logistic Regression Voice Balanced Accuracy 0.653 (0.551, 0.750) 0.663 (0.555, 0.761) 0.666 (0.563, 0.767) 0.619 (0.541, 0.689) 0.565 (0.487, 0.644) 0.528 (0.438, 0.603) Sensitivity 0.560 (0.360, 0.762) 0.640 (0.429, 0.833) 0.600 (0.400, 0.807) 0.711 (0.558, 0.849) 0.579 (0.428, 0.739) 0.526 (0.355, 0.676) Specificity 0.746 (0.713, 0.779) 0.685 (0.649, 0.721) 0.732 (0.698, 0.766) 0.528 (0.502, 0.557) 0.551 (0.524, 0.581) 0.530 (0.503, 0.555) AUC 0.724 (0.604, 0.834) 0.714 (0.637, 0.790) 0.719 (0.612, 0.825) 0.665 (0.578, 0.747) 0.603 (0.497, 0.704) 0.545 (0.439, 0.640) Voice + Demographics Balanced Accuracy 0.657 (0.544, 0.756) 0.797 (0.722, 0.859) 0.723 (0.623, 0.816) 0.661 (0.586, 0.735) 0.747 (0.683, 0.800) 0.701 (0.626, 0.774) Sensitivity 0.560 (0.345, 0.750) 0.880 (0.733, 1.000) 0.680 (0.500, 0.864) 0.711 (0.558, 0.849) 0.842 (0.714, 0.947) 0.684 (0.537, 0.829) Specificity 0.754 (0.722, 0.787) 0.715 (0.677, 0.749) 0.767 (0.732, 0.799) 0.612 (0.587, 0.639) 0.652 (0.627, 0.678) 0.718 (0.694, 0.743) AUC 0.733 (0.627, 0.827) 0.836 (0.767, 0.892) 0.813 (0.738, 0.883) 0.687 (0.607, 0.762) 0.777 (0.719, 0.828) 0.782 (0.721, 0.838)
Table 4: Classification results for each of the three acoustic features used alone and when combined with the demographics. These models are evaluated on both a holdout test set from the FEMH dataset and externally on the SVD dataset. Bold values are the highest for that metric and test set. SVM - support vector machine, MLP - multi-layered perceptron
The FEMH dataset includes data on 26 symptoms for each patient, which is not available in the SVD dataset. Consequently, classifiers incorporating symptom data cannot be evaluated on the external SVD dataset. Table 5 shows the results on the FEMH holdout test set when symptoms data are included as input to the classifier. Using this symptom data improves the classification performance across all algorithms and features. The best overall performance is found using logistic regression with symptom and demographic data combined with OpenSmile features, yielding a balanced accuracy of 83.7%.
We see an increase in accuracy when symptom data are included as input to the system. While this extra information is valuable, it is not easy to obtain. Although most clinics will obtain some medical history from their patients, this will vary between clinics, making it hard to generalise the system. There is also a concern that, should this be incorporated into an app for patient use, too many questions about their symptoms and medical history may be offputting and lead to inaccurate data entry as patients attempt to speed up the process.
Classifier Inputs Metric FEMH (holdout test dataset)
FeatureStates OpenSmile MFCC
SVM Voice + Demographics + Symptoms Balanced Accuracy 0.766 (0.671, 0.848) 0.807 (0.720, 0.872) 0.793 (0.707, 0.862)
Sensitivity 0.720 (0.526, 0.880) 0.840 (0.678, 0.962) 0.840 (0.667, 0.966)
Specificity 0.811 (0.780, 0.841) 0.775 (0.741, 0.807) 0.746 (0.714, 0.780)
AUROC 0.863 (0.799, 0.918) 0.893 (0.835, 0.937) 0.848 (0.796, 0.892)
Voice + Symptoms Balanced Accuracy 0.813 (0.720, 0.886) 0.814 (0.725, 0.891) 0.697 (0.597, 0.795)
Sensitivity 0.800 (0.609, 0.950) 0.800 (0.619, 0.950) 0.640 (0.438, 0.826)
Specificity 0.827 (0.796, 0.857) 0.828 (0.799, 0.856) 0.754 (0.722, 0.788)
AUROC 0.859 (0.785, 0.918) 0.866 (0.798, 0.918) 0.804 (0.723, 0.879)
MLP Voice + Demographics + Symptoms Balanced Accuracy 0.695 (0.596, 0.800) 0.776 (0.673, 0.859) 0.779 (0.676, 0.872)
Sensitivity 0.440 (0.241, 0.647) 0.720 (0.516, 0.885) 0.680 (0.483, 0.867)
Specificity 0.950 (0.931, 0.967) 0.831 (0.801, 0.859) 0.877 (0.852, 0.903)
AUROC 0.865 (0.787, 0.933) 0.892 (0.839, 0.937) 0.883 (0.828, 0.936)
Voice + Symptoms Balanced Accuracy 0.742 (0.641, 0.838) 0.792 (0.701, 0.873) 0.786 (0.698, 0.857)
Sensitivity 0.640 (0.440, 0.826) 0.760 (0.571, 0.920) 0.800 (0.619, 0.946)
Specificity 0.844 (0.817, 0.873) 0.824 (0.791, 0.851) 0.772 (0.739, 0.803)
AUROC 0.839 (0.764, 0.903) 0.889 (0.834, 0.936) 0.865 (0.802, 0.919)
Logistic Regression Voice + Demographics + Symptoms Balanced Accuracy 0.798 (0.711, 0.871) 0.837 (0.752, 0.901) 0.827 (0.754, 0.887)
Sensitivity 0.800 (0.633, 0.947) 0.840 (0.680, 0.966) 0.880 (0.737, 1.000)
Specificity 0.797 (0.765, 0.826) 0.833 (0.802, 0.862) 0.773 (0.741, 0.806)
AUROC 0.896 (0.844, 0.941) 0.918 (0.871, 0.954) 0.910 (0.863, 0.948)
Voice + Symptoms Balanced Accuracy 0.807 (0.723, 0.874) 0.786 (0.695, 0.869) 0.806 (0.719, 0.874)
Sensitivity 0.840 (0.677, 0.966) 0.760 (0.579, 0.923) 0.840 (0.667, 0.966)
Specificity 0.773 (0.739, 0.806) 0.813 (0.782, 0.842) 0.772 (0.738, 0.806)
AUROC 0.868 (0.797, 0.927) 0.879 (0.831, 0.922) 0.878 (0.821, 0.925)
Table 5: Classification results for the three acoustic features combined with the demographics and symptoms. These models could only be evaluated on a holdout test set from the FEMH dataset, as the external test set (SVD) does not include the required symptom data. Bold values are the highest for that metric and test set.
4.2 Fair Classification

### v2

@@ 3.4.1 Predictive Performance
In addition to the above metrics, we also suggest the Area Under the Receiver Operating Characteristic curve (AUROC). The ROC curve plots the true positive rate against the false positive rate at different decision thresholds. Since the ROC curve evaluates the model’s performance across multiple decision thresholds, the AUROC gives a measure of how good the model is regardless of the threshold chosen. This metric is also robust against imbalanced datasets as it does not bias towards the majority class.
@@ 3.4.2 Fairness
In medical applications, it is essential to assess the fairness of the algorithms being developed. As such, we suggest conducting statistical tests to evaluate the relationship between classifier performance and patient demographics. The most commonly available patient demographic data are sex and age of patients. For sex, we propose using a Fisher’s Exact Test to examine any association between classifier performance and patient sex Sprent (2011). We propose a Fisher’s Exact test as it is used for categorical variables and widely recommended for small sample sizes Kim (2016). For age, we suggest a t-test to compare performance across different age groups Kalpić et al. (2011).
@@ 4 Results and Discussion
In this section we present the results of the classification models on both the holdout FEMH test set and external SVD test set. As stated above we present the model’s predictive performances, fairness, and prediction times.
All experiments were performed using Python 3.10.12 on a machine running Windows 11 with an Intel Core i7-1260P CPU with 16GB RAM. All code and models are available on github at github.com/mary-paterson/LaryngealCancerClassificationBenchmark
@@ 4.1 Predictive Performance
First we present the predictive performance of the created models using the metrics described above. We compare the audio feature sets and classification algorithms as well as investigating how incorporating demographic and symptom data effects model performance.
4.1.1 Models using Audio Features Only
We start by discussing the results of the models that take the audio features only as input, these results can be seen in Table 4. These results are shown for both the FEMH (holdout) and SVD (external) test sets. Within the table the green highlighted cells show the results for the best models, while the best audio feature for each of the classification algorithms is highlighted in yellow. We also present the 95% confidence intervals for each result.
Algorithm Feature Set Test Set Balanced Acc. Sensitivity Specificity AUROC
FEMH
0.691
(0.593, 0.784)
0.680
(0.480, 0.857)
0.702
(0.665, 0.736)
0.760
(0.665, 0.845)
FeatureStates SVD
0.628
(0.557, 0.696)
0.763
(0.621, 0.900)
0.493
(0.466, 0.521)
0.649
(0.562, 0.735)
FEMH
0.680
(0.580, 0.772)
0.640
(0.440, 0.818)
0.720
(0.686, 0.754)
0.742
(0.676, 0.805)
OpenSMILE SVD
0.582
(0.497, 0.662)
0.500
(0.341, 0.667)
0.663
(0.637, 0.690)
0.615
(0.522, 0.711)
FEMH
0.631
(0.531, 0.731)
0.400
(0.206, 0.609)
0.861
(0.835, 0.887)
0.701
(0.594, 0.807)
SVM MFCC SVD
0.532
(0.462, 0.613)
0.289
(0.154, 0.455)
0.775
(0.752, 0.799)
0.592
(0.503, 0.679)
FEMH
0.683
(0.576, 0.784)
0.520
(0.318, 0.714)
0.846
(0.817, 0.873)
0.754
(0.643, 0.857)
FeatureStates SVD
0.628
(0.547, 0.705)
0.526
(0.368, 0.676)
0.729
(0.706, 0.753)
0.692
(0.607, 0.772)
FEMH
0.632
(0.533, 0.727)
0.560
(0.360, 0.750)
0.704
(0.668, 0.740)
0.709
(0.637, 0.779)
OpenSMILE SVD
0.610
(0.529, 0.689)
0.579
(0.417, 0.735)
0.640
(0.615, 0.667)
0.633
(0.538, 0.727)
FEMH
0.581
(0.478, 0.688)
0.480
(0.280, 0.697)
0.682
(0.646, 0.718)
0.638
(0.514, 0.746)
MLP MFCC SVD
0.548
(0.477, 0.616)
0.763
(0.622, 0.889)
0.333
(0.308, 0.358)
0.575
(0.489, 0.657)
FEMH
0.653
(0.551, 0.750)
0.560
(0.360, 0.762)
0.746
(0.713, 0.779)
0.724
(0.604, 0.834)
FeatureStates SVD
0.619
(0.541, 0.689)
0.711
(0.558, 0.849)
0.528
(0.502, 0.557)
0.665
(0.578, 0.747)
FEMH
0.663
(0.555, 0.761)
0.640
(0.429, 0.833)
0.685
(0.649, 0.721)
0.714
(0.637, 0.790)
OpenSMILE SVD
0.565
(0.487, 0.644)
0.579
(0.428, 0.739)
0.551
(0.524, 0.581)
0.603
(0.497, 0.704)
FEMH
0.666
(0.563, 0.767)
0.600
(0.400, 0.807)
0.732
(0.698, 0.766)
0.719
(0.612, 0.825)
Logistic Regression MFCC SVD
0.528
(0.438, 0.603)
0.526
(0.355, 0.676)
0.530
(0.503, 0.555)
0.545
(0.439, 0.640)
Table 4: The classification results for the models using audio features only as input to the system. The results of the best model are highlighted in green with the best audio feature for each of the classification algorithms highlighted in yellow. The confidence intervals are presented in brackets.
On the holdout test set, when using audio features only, FeatureStates performs best for both the SVM and MLP whereas, for logistic regression, MFCC performs the best. On the external test set, FeatureStates performs best for all three algorithms.
The best model using audio features alone as input is the SVM using FeatureStates. This model achieves a balanced accuracy of 69.1%, sensitivity of 68.0%, specificity of 70.2%, and AUROC of 76.0% on the FEMH holdout test set. On the external SVD test set a balanced accuracy of 62.8%, sensitivity of 76.3%, specificity of 49.3%, and AUROC of 64.9% is achieved.
These results show that there is no one audio feature set that works best across all algorithms, nor is there a single algorithm that has the best performance. The performance varies across the combinations of audio features and classification algorithms. This highlights the importance of a systematic approach to developing AI systems in terms of choosing both audio features and classification algorithms.
4.1.2 Incorporating Demographic Data
We now look to the results of the models that incorporate demographic data alongside audio features. The results of these models can be seen in Table 5. When comparing these results to those of the models that use audio features alone we see an improvement for all models.
Algorithm Feature Set Test Set Balanced Acc. Sensitivity Specificity AUROC
SVM FeatureStates FEMH
0.700
(0.594, 0.789
0.640
(0.429, 0.818
0.759
(0.726, 0.793
0.773
(0.682, 0.855
SVD
0.609
(0.529, 0.691
0.526
(0.364, 0.688
0.691
(0.666, 0.717
0.716
(0.642, 0.786
FEMH
0.752
(0.656, 0.837
0.760
(0.571, 0.923
0.743
(0.709, 0.776
0.802
(0.743, 0.851
OpenSMILE SVD
0.646
(0.566, 0.730
0.526
(0.370, 0.688
0.765
(0.741, 0.787
0.772
(0.714, 0.829
FEMH
0.701
(0.599, 0.803
0.600
(0.400, 0.800
0.802
(0.770, 0.832
0.807
(0.720, 0.881
MFCC SVD
0.681
(0.599, 0.759
0.605
(0.442, 0.756
0.758
(0.735, 0.781
0.769
(0.699, 0.831
FEMH
0.629
(0.522, 0.730
0.440
(0.227, 0.640
0.819
(0.786, 0.847
0.758
(0.663, 0.833
FeatureStates SVD
0.669
(0.586, 0.748
0.553
(0.390, 0.704
0.786
(0.764, 0.808
0.715
(0.627, 0.789
FEMH
0.755
(0.659, 0.841
0.760
(0.571, 0.923
0.750
(0.717, 0.784
0.809
(0.753, 0.858
OpenSMILE SVD
0.696
(0.612, 0.775
0.658
(0.488, 0.813
0.734
(0.708, 0.758
0.758
(0.695, 0.822
FEMH
0.655
(0.556, 0.762
0.440
(0.250, 0.654
0.871
(0.844, 0.897
0.805
(0.725, 0.881
MLP MFCC SVD
0.657
(0.573, 0.742
0.500
(0.333, 0.667
0.815
(0.794, 0.835
0.762
(0.695, 0.826
FEMH
0.657
(0.544, 0.756
0.560
(0.345, 0.750
0.754
(0.722, 0.787
0.733
(0.627, 0.827
FeatureStates SVD
0.661
(0.586, 0.735
0.711
(0.558, 0.849
0.612
(0.587, 0.639
0.687
(0.607, 0.762
FEMH
0.797
(0.722, 0.859
0.880
(0.733, 1.000
0.715
(0.677, 0.749
0.836
(0.767, 0.892
OpenSMILE SVD
0.747
(0.683, 0.800
0.842
(0.714, 0.947
0.652
(0.627, 0.678
0.777
(0.719, 0.828
FEMH
0.723
(0.623, 0.816
0.680
(0.500, 0.864
0.767
(0.732, 0.799
0.813
(0.738, 0.883
Logistic Regression MFCC SVD
0.701
(0.626, 0.774
0.684
(0.537, 0.829
0.718
(0.694, 0.743
0.782
(0.721, 0.838
Table 5: The classification results for the models using audio features and demographic data as input to the system. The results of the best model are highlighted in green with the best audio feature for each of the classification algorithms highlighted in yellow. The confidence intervals are presented in brackets.
OpenSMILE performs best across all three algorithms for the holdout test set. However, on the external test set, while OpenSMILE performed best for MLP and logistic regression, MFCC was best for SVM.
@@ 4.1.2 Incorporating Demographic Data
The best model using audio features and demographic data is the logistic regression with OpenSMILE features. This achieves a balanced accuracy of 79.7%, sensitivity of 88.0%, specificity of 71.5%, and AUROC of 83.6% on the FEMH holdout test set. On the external SVD test set a balanced accuracy of 74.7%, sensitivity of 84.2%, specificity of 65.2%, and AUROC of 77.7% is achieved.
When comparing the best model using audio features only and the best model using audio features and demographic data, an increase of 10.6% balanced accuracy is seen when demographic data is incorporated. This improvement is unsurprising due to the significantly different demographics seen between the classes for the two datasets. This difference in demographics between the benign and malignant patients is representative of what would be expected in real-world applications. This may however, cause problems with regards to the fairness of the models as the classifiers may have an over-reliance on demographic data. This would mean that any patient outside of the demographic majority (i.e. male over the age of 50) would not be classified as having a malignant pathology regardless of their voice recording.
In order to understand the robustness of the classifiers we can compare the performance of the models on the holdout FEMH test set and external SVD test set. Figure 5 shows the difference in balanced accuracy between the holdout FEMH and external SVD test sets. Although the performance on the external dataset is generally worse, the confidence intervals overlap for all classifiers except for the logistic regression using MFCC features without demographic data. The results also show that while the performance on the external test set is reduced, the model does not collapse on the external test set with a good balanced of sensitivity and specificity maintained.
Figure 5: The balanced accuracy for the holdout (FEMH) and external (SVD) test sets. 95% confidence intervals are shown.
Since Wav2Vec2 was trained on multiple datasets and is designed to be generalisable across different recording environments and recording devices, we expected this feature would be best suited to perform well on both the holdout and external test sets. This is seen with the models using audio features only with the FeatureStates models having the smallest decrease in performance between the FEMH and SVD test sets. This is not seen for the models that use audio features and demographic data, possibly since the demographic data dominates the classification decision.
4.1.3 Incorporating Symptom Data
In addition to the audio recordings and demographic data, the FEMH dataset also includes 26 data points for each patient their symptoms. We incorporate this symptom data into classifiers to investigate whether this may improve classifier performance. This data is unavailable in the SVD, so these classifiers cannot be tested on this external test set.
We start by incorporating the symptom data with the audio features without using the demographic data, the results of these models can be seen in Table 6. The OpenSMILE features perform best for both the SVM and MLP, with the MFCCs performing best for the logistic regression. The best performing classifier is the SVM with OpenSMILE features achieving a balanced accuracy of 81.4%, sensitivity of 80.0%, specificity of 82.8%, and AUROC of 86.6%. For the majority of models the use of symptom data instead of demographic data improves model performance. There are only two exceptions: the SVM using MFCCs and the logistic regression using OpenSMILE features. This suggests that the symptom data is more valuable than the demographic data.
Algorithm Feature Set Balanced Acc. Sensitivity Specificity AUC
FeatureStates
0.813
(0.720, 0.886
0.800
(0.609, 0.950
0.827
(0.796, 0.857
0.859
(0.785, 0.918
OpenSMILE
0.814
(0.725, 0.891
0.800
(0.619, 0.950
0.828
(0.799, 0.856
0.866
(0.798, 0.918
SVM MFCC
0.697
(0.597, 0.795
0.640
(0.438, 0.826
0.754
(0.722, 0.788
0.804
(0.723, 0.879
FeatureStates
0.742
(0.641, 0.838
0.640
(0.440, 0.826
0.844
(0.817, 0.873
0.839
(0.764, 0.903
OpenSMILE
0.792
(0.701, 0.873
0.760
(0.571, 0.920
0.824
(0.791, 0.851
0.889
(0.834, 0.936
MLP MFCC
0.786
(0.698, 0.857
0.800
(0.619, 0.946
0.772
(0.739, 0.803
0.865
(0.802, 0.919
FeatureStates
0.807
(0.723, 0.874
0.840
(0.677, 0.966
0.773
(0.739, 0.806
0.868
(0.797, 0.927
OpenSMILE
0.786
(0.695, 0.869
0.760
(0.579, 0.923
0.813
(0.782, 0.842
0.879
(0.831, 0.922
Logistic Regression MFCC
0.806
(0.719, 0.874
0.840
(0.667, 0.966
0.772
(0.738, 0.806
0.878
(0.821, 0.925
Table 6: The classification results for the models using audio features and symptom data as input to the system. The results of the best model are highlighted in green with the best audio feature for each of the classification algorithms highlighted in yellow. The confidence intervals are presented in brackets.
Next, we combine the audio features, demographic, and symptom data. This combination yields the best classification results. Table 7 shows the results of these classifiers. The OpenSMILE features perform best for all three algorithms. The best model is logistic regression using OpenSMILE features which achieves a balanced accuracy of 83.7%, sensitivity of 84.0%, specificity of 83.3%, and AUROC of 91.8%.
Algorithm Feature Set Balanced Acc. Sensitivity Specificity AUROC
FeatureStates
0.766
(0.671, 0.848)
0.720
(0.526, 0.880)
0.811
(0.780, 0.841)
0.863
(0.799, 0.918)
OpenSMILE
0.807
(0.720, 0.872)
0.840
(0.678, 0.962)
0.775
(0.741, 0.807)
0.893
(0.835, 0.937)
SVM MFCC
0.793
(0.707, 0.862)
0.840
(0.667, 0.966)
0.746
(0.714, 0.780)
0.848
(0.796, 0.892)
FeatureStates
0.695
(0.596, 0.800)
0.440
(0.241, 0.647)
0.950
(0.931, 0.967)
0.865
(0.787, 0.933)
OpenSMILE
0.776
(0.673, 0.859)
0.720
(0.516, 0.885)
0.831
(0.801, 0.859)
0.892
(0.839, 0.937)
MLP MFCC
0.779
(0.676, 0.872)
0.680
(0.483, 0.867)
0.877
(0.852, 0.903)
0.883
(0.828, 0.936)
FeatureStates
0.798
(0.711, 0.871)
0.800
(0.633, 0.947)
0.797
(0.765, 0.826)
0.896
(0.844, 0.941)
OpenSMILE
0.837
(0.752, 0.901)
0.840
(0.680, 0.966)
0.833
(0.802, 0.862)
0.918
(0.871, 0.954)
Logistic Regression MFCC
0.827
(0.754, 0.887)
0.880
(0.737, 1.000)
0.773
(0.741, 0.806)
0.910
(0.863, 0.948)
Table 7: The classification results for the models using audio features, demographic, and symptom data as input to the system. The results of the best model are highlighted in green with the best audio feature for each of the classification algorithms highlighted in yellow. The confidence intervals are presented in brackets.
Although all of these models improve upon the models using audio features only and the models using audio features and demographic data, the addition of demographic data does not improve all of the models compared to incorporating simply the symptom data. In fact, most models that use audio features and symptom data are not improved by the addition of demographic data. The three models that are improved by the incorporation of demographic data are the SVM and logistic regression models using MFCCs, and the logistic regression model using OpenSMILE features. This suggests that the use of demographic data is not required to improve classification performance should symptom data be available. This may make models fairer as classification would not be able to be made directly on the basis of the patient’s age or sex.
The incorporation of symptom data as an input to the classifier improves classification performance, however, this data is not simple to obtain. Although most clinics will obtain some medical history from their patients, this will vary between clinics, making it hard to generalise the system since external validation requires the same symptoms to have been collected. Additionally, the same symptoms would need to be collected should future researchers want to extend this work by collecting more data. There is also a concern that data collection be facilitated by an app for patient use, too many questions about their symptoms and medical history may be off-putting and risks being inaccurate if patients were to get frustrated with a long process.
4.1.4 Comparing Our Work with Others
To understand how well our methods perform, we can compare the classifier performance to those reported by Wang et al. Wang et al. (2024). Wang et al. also perform the same classification task as this work as well as using the same datasets (FEMH for training and testing, SVD for external validation). While there are many articles published within this area of research, Wang et al. present the only article using the FEMH dataset for the classification of benign and malignant pathologies meaning that their results can be directly compared to those we present. This allows us to understand how the methods presented in this paper compare to other work in this field. Wang et al. use a deep neural network (DNN) to classify patients, using the whole voice signal as input.
It is important to note some differences in the article presented by Wang et al. compared to this work. Wang et al. reports accuracy as opposed to balanced accuracy. As such, we calculate balanced accuracy for Wang et al. by averaging the reported sensitivity and specificity values. Balanced accuracy is a much more meaningful metric in this task due to the highly imbalanced dataset. Wang et al. did not test on a holdout set of the FEMH data but reported average results from 10-fold cross-validation. Wang et al. reports a different number of patients in the malignant and benign classes in the SVD, possibly due to our work including pre-malignant pathologies and limiting patients to those over 18.
Unfortunately, Wang et al. do not state what pathologies they included in the benign and malignant groups from the SVD, nor do they state that there were any exclusion criteria applied to the patients in the SVD. Wang et al. also do not create a model using audio features and symptom data without demographic data, as such we cannot compare the results of those models to Wang et al..
While we are able to compare our work to that of Wang et al. these differences highlight the need for standardization and open-science. When open-source datasets and code are not shared it becomes impossible to meaningfully compare work even when they perform the same task. Even when models are trained on private datasets, if they are tested on publicly available datasets their results can be compared to other models. This can also be done by sharing the trained models which would allow them to be tested on new datasets whenever they are made available. By standardizing evaluation methods we can also increase the comparability of models. When work only presents a single metric it can be impossible to meaningfully compare their results to those of other work. This is especially pertinent when accuracy is the only reported result despite the use of highly imbalanced test sets.
Despite the difficulties in comparing work in this area we feel that we can meaningfully compare our results to those reported by Wang et al.. In Table 8 we compare the results of the best models for each combination of input variables from our work to those reported by Wang et al. the best result for each metric for each of the two datasets is bolded.
FEMH (holdout test dataset) SVD (external test dataset)
Model Input Metric Wang et al. Wang et al. (2024) Our Results Wang et al. Wang et al. (2024) Our Results
Voice Balanced Accuracy 0.545 0.691 0.455 0.628
Sensitivity 0.550 0.680 0.480 0.763
Specificity 0.540 0.702 0.430 0.493
AUROC 0.631 0.760 0.470 0.649
Voice + Demographics Balanced Accuracy 0.690 0.797 0.742 0.747
Sensitivity 0.680 0.880 0.739 0.842
Specificity 0.700 0.715 0.745 0.652
AUROC 0.807 0.836 0.785 0.777
Voice + Demographics + Symptoms Balanced Accuracy 0.800 0.837 N/A N/A
Sensitivity 0.783 0.840 N/A N/A
Specificity 0.816 0.833 N/A N/A
AUROC 0.878 0.918 N/A N/A
Table 8: Results obtained in this work compared to results reported by Wang et al. (2024). Bolded values are the best within that classifier type. For our work the classifier with the highest balanced accuracy was chosen for each input type. For Voice our results are from the SVM with FeatureStates input, for Voice + Demographics our results are from the logistic regression with OpenSMILE input, and for the Voice + Demographics + Symptoms our results are from the logistic regression with OpenSMILE input.
When using the audio as the only input to the system, we outperform the results reported by Wang et al. in all metrics for both the FEMH and SVD test sets. We improve on balanced accuracy by 14.6%, sensitivity by 13.0%, specificity by 16.2%, and AUROC by 12.9% for the FEMH dataset. For the external SVD test set we improve balanced accuracy by 17.3%, sensitivity by 28.3%, specificity by 6.3%, and AUROC by 17.9%.
When demographics are incorporated, we outperform Wang et al. for all metrics on the FEMH test set, however, on the external test set, we only perform better in balanced accuracy and sensitivity. For the FEMH test set we outperform Wang et al.’s balanced accuracy by 10.7%, sensitivity by 20.0%, specificity by 1.5%, and AUROC by 2.9%. For the external SVD test set differences are marginal with only an increase of 0.5% in balanced accuracy, and while a significant increase of 10.3% is seen for the sensitivity this is offset by the 9.3% reduction in specificity. A small reduction of 0.8% is seen for the AUROC.
When symptoms and demographics are incorporated we outperform Wang et al. in all metrics on the FEMH test set (as previously stated these models cannot be tested on the external dataset as it lacks matching symptom data). With an increase of 3.7% for balanced accuracy, 5.7% for sensitivity, 1.7% for specificity, and 4.0% for AUROC.
These results suggest that, although the AI methods employed in this work are more simplistic than those used by Wang et al. they perform better. The results for the audio only model suggest that this model is more generalizable across different datasets than the method presented by Wang et al. since our results are significantly better on the external SVD dataset (increase in balanced accuracy of 17.3%).
4.2 Fairness